OpenStudy (anonymous):

I WILL FAN AND MEDAL!!! What is the value of the x variable in the solution to the following system of equations? 3x − 6y = 3 7x − 5y = −11 −2 −3 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system.

OpenStudy (anonymous):

the answer is -2, let me explain why

HanAkoSolo (jamierox4ev3r):

http://openstudy.com/code-of-conduct @AwesomeB plz show process before then give answer, not the other way around. Thank you! ^.^

OpenStudy (anonymous):

I know the proccess I just wanted to see if the answer I had was right.

OpenStudy (anonymous):

multiply 3x - 6y = 3 by -7: -7(3x - 6y) = -7(3) -21x + 42y = -21 and then i will multiply 7x - 5y = -11 by 3: 21x - 15y = -33 add them: 21x - 21x - 15y + 42y = -21 - 33 27y = -54 y = -2

OpenStudy (anonymous):

Thanks! @AwesomeB

OpenStudy (anonymous):

the answer is -3.

OpenStudy (anonymous):

I'm confused now...

OpenStudy (anonymous):

yw and @ranjeet619 the correct answer is -2 check my work

HanAkoSolo (jamierox4ev3r):

wait a second she is asking the value of the x-variable, not the y-variable. SO you are incorrect @AwesomeB

HanAkoSolo (jamierox4ev3r):

the y-variable is correct, but the x-variable is undefinded as of now. @ranjeet may be correct, but i will prove for sure

HanAkoSolo (jamierox4ev3r):

3x-6(-2)=3 3x+12=3 -12 -12 3x=-9 /3 /3 x=-3 7x-5(-2)=-11 7x+10=-11 -10 -10 7x=-21 /7 /7 x=-3

HanAkoSolo (jamierox4ev3r):

therefore, x=-3

OpenStudy (anonymous):

the coordinates are (-3-2) @ranjeet619 was correct my fault

HanAkoSolo (jamierox4ev3r):

does that make sense @aleximhappy08 ?

OpenStudy (anonymous):

(-3,-2)

HanAkoSolo (jamierox4ev3r):

right but the question is asking what the x-variable is, and that is -3. That is why you MUST show your work before you provide an answer! I hope you learned your lesson @AwesomeB

OpenStudy (anonymous):

I read the question wrong and I thought she was looking for y

OpenStudy (anonymous):

Thanks guys!

HanAkoSolo (jamierox4ev3r):

anytime just hit me up if u need help in the future @aleximhappy08